OpenStudy (i_ask_alot):

A jar contains 11 pennies and 8 dimes. What is the ratio of dimes to pennies?

OpenStudy (i_ask_alot):

@rebeccaxhawaii

OpenStudy (anonymous):

im not 100% sure, but i think its 8 over 11

OpenStudy (anonymous):

1:1.375, for every dime there is 1.375 pennies

OpenStudy (bookworm14):

A ratio is a comparison they are represented as a:b (a and b are your coefficients aka numbers) in this case your answer would be 11:8 but if they want it simplified you do a/b or 11/8 = 1 : 1.375

OpenStudy (i_ask_alot):

Thank you @Bookworm14 @porkbuns @goopyfish908

OpenStudy (bookworm14):

Your welcome love! :)

OpenStudy (i_ask_alot):

Do you mind if I ask about 4 more? @Bookworm14

OpenStudy (bookworm14):

i don't mind! but with me you must participate ok? like if i ask you, give me your best guess if you dont know :) i would love to help you, but make sure you post them in new questions so u dnt jam up this one ok? :)

OpenStudy (i_ask_alot):

OK. This is the next one I think it is C A vase contains 12 tulips and 20 roses. What is the ratio of roses to the total number of flowers in the vase?

OpenStudy (i_ask_alot):

A. 3 to 8 B. 3 to 5 C. 5 to 8 D. 5 to 3

OpenStudy (bookworm14):

not quite, but your close! so we have to find the gcf (greatest common factor) write out your factors for 12 its (1,2,3,4,6,12) for 20 its (1,2,4,5,10,20) so our greatest factor is 4 divide them both by 4 12/4 = 3 20/4 = 5 our new simplified ratio is 3:5 !!

OpenStudy (i_ask_alot):

A recreational center has 15 basketballs and 12 soccer balls. What is the ratio of soccer balls to basketballs in simplest form? Is it 5:4

OpenStudy (bookworm14):

Yes it is! :D

OpenStudy (i_ask_alot):

Sylvia has a total of 16 trading cards: 8 superhero cards, 4 princess cards, 2 villain cards, and 2 wildlife cards. What is the ratio of wildlife cards to the total number of trading cards in simplest form? Is it 1:2

OpenStudy (bookworm14):

close, but not exactly, heres how: she has a total of 16 trading cards , she only has 2 wildlife cards (pay close attention though, they want the ratio of wildlife to trading, they are very specific lol) so this would be 2:16 or in simplest form 1:8 :)

OpenStudy (i_ask_alot):

Oh, OK I get it now that one was tricky this last one is the same way.

OpenStudy (i_ask_alot):

A bake sale has 12 cookies, 9 pies, 8 cakes, 4 tarts, and 3 cobblers to sell. What is the ratio of the total number of baked goods to cookies in simplest form? Is it 4:3

OpenStudy (bookworm14):

for this one there are 36 total baked goods. there are only 12 cookies this would be 36:12 36:12 reduces to 3:1

OpenStudy (bookworm14):

oh shoot, i have to go, im sorry!

OpenStudy (i_ask_alot):

OK Thanks for all your help!!!